# TypedThinker: Typed Thinking Improves Large Language Model Reasoning

### v1

@@ Abstract
Despite significant advancements in the reasoning capabilities of Large Language Models (LLMs), the lack of diverse reasoning solutions often makes them trapped in a limited solution search area. In this paper, we propose TypedThinker, a reasoning framework that diversifies LLMs’ reasoning solutions by incorporating multiple reasoning types (deductive, inductive, abductive, and analogical). Our analysis across four benchmarks reveals that different reasoning types uniquely solve distinct sets of problems, highlighting the importance of diverse thinking approaches. TypedThinker addresses two key challenges: selecting appropriate reasoning types for given problems and effectively implementing specific reasoning types. Through self-training on successful experiences, TypedThinker learns an implicit policy for reasoning type selection and application. Experimental results demonstrate significant improvements over baseline models, with accuracy increases of 3.4% for Mistral 7B and 16.7% for LLaMA3 8B across four reasoning benchmarks. Notably, TypedThinker shows effective generalization to new benchmarks and can further enhance the reasoning capability of powerful models like GPT-4o. The code is released at https://github.com/dqwang122/ThinkHub.
@@ 1 Introduction
Large Language Models (LLMs) exhibited promising capabilities in reasoning, such as solving logical reasoning and mathematical problems (Bai et al., 2022; OpenAI, 2023). Plenty of work has been done to improve the reasoning capabilities by adding reasoning thoughts (Wei et al., 2022) and making these thoughts more elaborated (Fu et al., 2023; Zheng et al., 2024). However, the exploration of novel reasoning thoughts is understudied. The lack of diversity in reasoning makes LLMs easily trapped in a fixed mindset, which limits their performance in solving difficult problems.
Current research efforts to enhance reasoning diversity remain unsatisfactory. AlphaCode (Li et al., 2022; Leblond et al., 2023) randomizes the difficulty level and categorical tags of the code problems in the prompt to encourage diversity. However, its dependency on manually curated attributes makes it challenging to scale and apply beyond coding problems. On the other hand, increasing the temperature is an easy way to generate superficially diverse outputs, but it often fails to produce high-quality reasoning solutions with significant differences. For example, repeated sampling (Brown et al., 2024) generates 100,000 solutions per problem with temperature 0.6, but their solutions  [^1] Their results can be found at: https://huggingface.co/datasets/ScalingIntelligence/monkey_business. are mostly based on deductive reasoning, which starts from the problem context to infer the answer step by step.
Encouraging LLMs to apply more types of reasoning can make them think more diversely. Besides deductive reasoning, there are other reasoning types such as inductive (Flach and Kakas, 2000), abductive (Douven, 2011), and analogical reasoning (Bartha, 2013). These reasoning types reflect different mental methods of drawing conclusions, leading to diverse thinking processes. This is also inspired by human cognitive processes, where individuals selectively apply contextually appropriate logical reasoning strategies (Halpern, 2014; Bronkhorst et al., 2020). For example, when some answer candidates are given for a free-response math problem, humans are more likely to verify the correctness of each option instead of solving it from the given conditions. This reflects human’s implicit mental shift in the type of reasoning from deductive to abductive.
To investigate how different reasoning types affect the LLMs’ performance in solving reasoning problems, we calculate the percentage of problems that can only be solved by one particular reasoning type in Figure 1. Specifically, we generate 10 samples for each reasoning type with the Mistral 7B instruct (Jiang et al., 2023) on four benchmarks: LogiQA (Liu et al., 2023a) and BBH (Suzgun et al., 2022), GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). We consider that one problem is solved by a reasoning type if at least one correct solution is found among the 10 samples of this reasoning type. For example, the blue region in Figure 1 reflects the percentage of problems that can only be solved by inductive reasoning, which is surprisingly effective on the MATH benchmark. It means that without inductive reasoning, LLMs can hardly find a correct solution for these 11.03% problems. As we can see, each reasoning type has a unique set of problems that can be solved by it, highlighting the importance of divergent thinking based on the reasoning types to enlarge the solvable problem set. More analysis can be found in Section  4.2.
Figure 1: The percentage distribution of each reasoning type for problems. The y-axis indicates the percentage of problems that can be only solved by one specific reasoning type.
Integrating diverse reasoning strategies into LLMs’ problem-solving capabilities has two main challenges. First, as shown in Figure 1, it is important to select the appropriate reasoning type for a given problem, otherwise, it may mislead the LLMs with the wrong thinking direction. Besides, it may be difficult for LLMs to follow a specific reasoning type to solve the problem. Therefore, we propose a novel approach TypedThinker with the implicit policy to select and conduct the appropriate reasoning types and the explicit memory to retrieve relevant experiences to aid problem-solving. TypedThinker uses the meta-thinker and the reasoner, to select the reasoning types and conduct the selected types of reasoning. It also maintains an explicit memory to retrieve relevant knowledge and experiences. To learn the implicit policy for reasoning types, TypedThinker is optimized based on its own successful experiences. For each problem in the training set, it estimates the effectiveness score of each reasoning type based on the success rate during sampling. The meta-thinker is fine-tuned on these empirical effectiveness scores to learn reasoning type selection for each problem. These successful reasoning experiences are also used to enhance the reasoner’s capability of following a specific reasoning type in problem-solving. During the self-training process, TypedThinker updates the implicit policy of the meta-thinker to select reasoning types and the policy of the reasoner to apply this type, while storing the successful reasoning experiences in the explicit memory for retrieval.
@@
Experimental results show that TypedThinker improves Mistral 7B instruct (Jiang et al., 2023) by 3.4% and LLaMA3 8B instruct (Touvron et al., 2023) by 16.7% on the average accuracy of two logical benchmarks, LogiQA and BBH, and two mathematics benchmarks, GSM8k and MATH. We further demonstrate that TypedThinker can directly be applied to the new benchmark Contexthub (Hua et al., 2024) and outperforms other baselines. Moreover, we show that the meta-thinker based on the weak 7B model can enhance the performance of GPT-4o without distilling knowledge from strong LLMs, which also sheds light on the weak-to-strong generalization (Burns et al., 2024).
@@ 2 Related Work
Logical Reasoning Logical reasoning includes various methods to emulate human-like thought processes (Wason and Johnson-Laird, 1972). Deductive reasoning focuses on deriving specific conclusions from general principles or premises, ensuring that conclusions logically follow if the premises are true (Johnson-Laird, 2010). In contrast, inductive reasoning involves generalizing from specific instances to broader principles, often used to identify patterns and make predictions based on empirical data (Flach and Kakas, 2000). Abductive reasoning, considered more creative and open-ended, involves forming hypotheses to explain observations, often generating the most plausible explanation rather than a guaranteed conclusion (Douven, 2011). Analogical reasoning is concerned with the comparison between two or more objects and drawing a conclusion based on the similarity (Bartha, 2013). Previous LLMs studies on logical reasoning mainly focus on benchmarking its performance in different reasoning types (Bang et al., 2023; Dougrez-Lewis et al., 2024; Luo et al., 2023; Yu et al., 2024). Instead, this paper mainly focuses on the selection and application of the appropriate reasoning type when solving a specific problem.
@@
Reasoning in Large Language Models Plenty of studies have been done to enhance the reasoning capability of LLMs. Chain-of-thoughts methods focus on creating better instructions to improve the quality of the reasoning process, such as Complex CoT (Du et al., 2023), Tree of Thought (ToT) (Yao et al., 2023) and Graph of Thought (Besta et al., 2024). Refinement-based methods revise LLMs solutions by the feedback from themselves or others model (Akyürek et al., 2023; Wang and Li, 2023; Du et al., 2023). Search-based methods use the reward model to search the best reasoning path (Lightman et al., ; Liu et al., 2023b; Hao et al., 2023). While most of them focus on creating high-quality reasoning paths, the diversity of thinking is understudied. Our paper aims to diversify the thinking process by incorporating different reasoning types.
@@
Self-improvement and Self-training in LLMs Recent works explore the self-improvement capability of LLMs, by finetuning LLMs on their high-quality generations (Wang et al., 2023; Huang et al., 2023; Toshniwal et al., 2024). This process can be extended to multiple iterations Gülçehre et al. (2023); Aksitov et al. . Benefiting the LLMs’ ability to follow instructions, researchers also ask LLMs to provide feedback themselves and improve their responses without finetuning (Peng et al., 2023; Shinn et al., 2023). This can be further enhanced by using their own feedback as the reward model to provide better signals for finetuning (Yuan et al., 2024; Kumar et al., 2024). In this paper, we focus on stimulating their capabilities to conduct various reasoning types and use these experiences to diversify their thinking in reasoning type selection and following.
@@ 3 TypedThinker: Diversify Thinking with Typed Reasoning
In this paper, we focus on four logical reasoning types: deductive, inductive, abductive, and analogical reasoning. For each reasoning type, we provide a short definition and a simple example to demonstrate the inference rules, which are listed in Table 8 in the Appendix. Based on that, we introduce a reasoning framework TypedThinker to diversify LLMs’ thinking with different reasoning types. As shown in Figure 2, there are three components in TypedThinker: the meta-thinker to select reasoning type, the finetuned reasoner to conduct specific reasoning, and explicit memory to retrieve experience. TypedThinker optimizes the implicit policy of the meta-thinker and reasoner and updates the explicit memory based on self-training.
Figure 2: TypedThinker consists of three components: the meta-thinker to select the reasoning types, the explicit memory to retrieve relevant experience, and the reasoner to conduct the specific reasoning. The meta-thinker is fine-tuned to predict an effective score $s\in[0,1]$ for each reasoning type.
@@
3.1 Typed Reasoning with the Implicit Policy and Explicit Memory
@@
Let $D=\{(\bm{x_{1}},y_{1}),\cdots,(\bm{x_{N}},y_{N})\}$ be a set of $N$ problems, where $\bm{x}_{i}$ and $y_{i}$ is the problem and the ground-truth answer of the $i-$th instance. We define a reasoning type space $\mathcal{F}$ that includes an empty type and four types of reasoning: deductive, inductive, abductive, and analogical. The goal is to model the selection and implementation of various reasoning types as implicit and explicit memory, thus enhancing LLMs’ performance in reasoning tasks. Here we use the exact match between the ground-truth answer $y$ and the system output $\hat{y}$ to calculate the accuracy as the measurement of task performance.
@@
Meta-thinker to select the reasoning type. Given a problem $\bm{x}$, the goal of the meta-thinker is to select an appropriate set of reasoning types to solve the problem. Specifically, it predicts an effectiveness score $s_{k}\in[0,1]$ for each reasoning type $f_{k}\in\mathcal{F}$, which can be represented as $s_{x,k}=\pi_{\theta}(\bm{x},f_{k})$. $s_{x,k}=0$ indicates that the problem $\bm{x}$ can hardly be solved by the reasoning type $f_{k}$ with a limited sampling times  [^2] In this paper, we sample at most 10 times for one problem.. Note that the effectiveness scores of different types are independent of each other and their sums are not necessary to be 1. The most effective reasoning type is defined as $f^{*}(\bm{x})=\arg\max_{f_{k}\in\mathcal{F}}s_{x,k}$. Meanwhile, we can obtain a set of reasoning types with a non-zero effective ratio, which we call the effective set: $F(\bm{x})=\{f_{k}|s_{x,k}>0\}$. We initialize $\pi_{\theta}$ with the pre-trained LLM and later fine-tune the parameters $\theta$ during the self-training. The prompt is listed in Appendix A.1.
@@
Explicit Memory to retrieve experience. TypedThinker keeps an explicit memory $M=\bigcup_{f_{k}\in\mathcal{F}}M_{k}$ for each type of reasoning. For each problem, we keep one correct solution per reasoning type if applicable, resulting in a set of at most $|D|\times|\mathcal{F}|$ solutions. If multiple solutions exist for one problem, we keep the longest to get a more detailed context. The entry in the memory $\mathcal{M}_{k}$ is represented as a tuple of $(\bm{x}_{r},sol_{r},)$. Here $sol_{r}$ is the concrete reasoning process of the reasoning type $f_{k}$, including the predicted answer. Given a new problem $\bm{x}$ and its reasoning type $f_{k}$, it retrieves a set of relevant experience $d_{x,k}=\{(\bm{x}_{r},sol_{r})\in M_{k}|L(\bm{x}_{r},\bm{x})<\delta\}$. $L$ is the distance function measuring relevancy between two problems and $\delta\in[0,1]$ is the relevancy threshold. In this study, we use the cosine similarity for the distance function. In this study, we use the cosine similarity between the semantic embeddings as the distance function. The retrieved experiences are used as the few-shot examples of the reasoner.
@@
Reasoner to perform the reasoning according to the type. The reasoner applies the reasoning type $f_{k}$ to the problem $\bm{x}$ and provides a detailed reasoning path for its predicted answer $\hat{y}$. The reasoner is based on LLM to conduct reasoning and the instruction is composed of $(\bm{x},f_{k},d_{x,k})$, where the $d_{k}$ is the retrieved relevant successful experience. The reasoner can be further optimized via instruction tuning to enhance the capability of conducting a specific type of reasoning.
@@
During inference, TypedThinker first uses the meta-thinker $\pi_{\theta}$ to predict an effective score $s_{k}$ for each reasoning type. It then retrieves relevant reasoning experience $d_{k}$ from the explicit memory. Finally, the reasoner conducts the specific type reasoning $f_{k}$ with the help of past successful experience. There are two approaches to aggregating multiple solutions. One is to greedily resample several times based on the most effective reasoning type $f^{*}$ and use self-consistency (Wang et al., 2022) to enhance the answer. The other is to sample solutions for all effective reasoning types $F$, and apply a weighted vote with the effective score as the coefficient. By default, we use the greedy approach for TypedThinker, and we discuss the weighted vote in Section 4.4.
3.2 Optimize Implicit Policy for Reasoning Type Selection and Following

### v2

# TypedThinker: Diversify Large Language Model Reasoning with Typed Thinking

@@ Abstract
Large Language Models (LLMs) have demonstrated strong reasoning capabilities in solving complex problems. However, current approaches primarily enhance reasoning through the elaboration of thoughts while neglecting the diversity of reasoning types. LLMs typically employ deductive reasoning, proceeding step-by-step from given conditions, which limits their exploration during problem-solving. Our analysis reveals that certain problems are exclusively solvable through specific reasoning strategies like inductive, abductive, or analogical reasoning. However, incorporating diverse reasoning approaches presents two key challenges: identifying the appropriate reasoning type for each problem and exploiting this approach during problem-solving. Therefore, we propose the TypedThinker that predicts suitable reasoning types based on the problem and their previous effectiveness and provides relevant demonstrations to guide LLMs in applying these strategies. Experimental results show significant improvements across multiple benchmarks, with performance gains of 3.4% for Mistral 7B, 6.5% for LLaMA3 8B, and 7% for Qwen 2 7B on logical and mathematical reasoning tasks. TypedThinker enhances LLM reasoning without requiring knowledge distillation from larger models. It can be integrated into more advanced systems like GPT-4o or specialized models like MetaMath to diversify their reasoning approaches and improve their problem-solving capabilities.
@@ 1 Introduction
Large Language Models (LLMs) exhibited promising capabilities in reasoning, such as solving logical reasoning and mathematical problems (Bai et al., 2022; OpenAI, 2023). Plenty of work has been done to improve the reasoning capabilities by adding reasoning thoughts (Wei et al., 2022) and making these thoughts more elaborated (Fu et al., 2023; Zheng et al., 2024). However, the exploration of diverse reasoning approaches remains severely limited. LLMs typically rely on a single reasoning pattern—usually deductive reasoning that proceeds step-by-step from given conditions. This narrow focus traps models in fixed thinking patterns, preventing them from solving problems that require different high-level reasoning approaches.
Current research fails to create truly diverse reasoning approaches. AlphaCode (Li et al., 2022; Leblond et al., 2023) tries to increase diversity by randomizing problem difficulty and tags. This method has limited scalability because it needs manual curation of attributes. It also works poorly beyond coding tasks. Increasing temperature settings offers another approach. This can generate outputs that appear different on the surface. Yet it rarely produces solutions with fundamentally different reasoning strategies. For example, repeated sampling (Brown et al., 2024) generated 100,000 solutions per problem using temperature 0.6. Their solutions  [^1] Their results are available at: https://huggingface.co/datasets/ScalingIntelligence/monkey_business. all follow the same basic approach. They start with problem conditions and work forward to deduce answers step-by-step. Humans, however, can use multiple reasoning strategies. One alternative approach is to propose a hypothesis first and then verify this hypothesis within the problem context. Current methods do not capture this diversity of thinking patterns.
Similarly, encouraging LLMs to explore diverse high-level thinking patterns leads to more varied solutions. Human cognitive research (Halpern, 2014; Bronkhorst et al., 2020) reveals multiple problem-solving approaches beyond deduction, including inductive (Flach and Kakas, 2000), abductive (Douven, 2011), and analogical reasoning (Bartha, 2013). These non-deductive strategies offer different directions through the solution space, often leading to more effective results. For example, abductive reasoning—proposing hypotheses and verifying them—is more suitable in multiple-choice scenarios. Humans demonstrate remarkable flexibility by switching from deductive to abductive reasoning when a free-response problem is reformatted with answer choices, despite unchanged problem content. This adaptive behavior shows how humans intuitively select optimal reasoning strategies for each problem format.
To examine how reasoning types affect LLM performance, we explicitly directed LLMs to apply specific reasoning strategies to problems. For each problem, the model is asked to follow a specific reasoning type and generate multiple solutions. We considered a problem "solvable" by a reasoning type if at least one of these solutions is correct. We calculate the percentage of problems solvable exclusively by one particular reasoning type. These represent cases where lacking the right reasoning approach makes problems nearly impossible to solve. We tested Mistral 7B instruct (Jiang et al., 2023) across four benchmarks: LogiQA (Liu et al., 2023a), BBH (Suzgun et al., 2022), GSM8k (Cobbe et al., 2021), and MATH (Hendrycks et al., 2021). Results in Figure 1 reveal that each reasoning type uniquely solves certain problems that other approaches cannot. This demonstrates that incorporating diverse reasoning strategies effectively expands the range of solvable problems. More detailed analysis is put in Section  4.2.
Figure 1: The percentage of problems solvable exclusively by one reasoning type.
However, incorporating reasoning types into LLM problem-solving faces two major challenges. First, identifying the right reasoning type is difficult. Figure 1 shows that incorrect reasoning types mislead LLMs with wrong thinking directions. Second, ensuring the model follows the chosen reasoning type during problem-solving is crucial. Therefore, we propose TypedThinker to address these challenges. It predicts suitable reasoning types based on previous successful experiences and uses explicit demonstrations to help LLMs effectively apply them. TypedThinker consists of three key components: the LLM reasoner, the meta-thinker and demonstration collection. The meta-thinker is fine-tuned on empirical effectiveness scores of each reasoning type from the training set. We use rejection sampling to collect successful solutions for each type as demonstrations, which enhance LLMs’ ability to exploit specific reasoning strategies. During inference, TypedThinker uses the meta-thinker to identify the most suitable reasoning type and retrieves relevant demonstrations to guide the LLM reasoner in applying this approach to solve the problem.
@@
Experimental results show that TypedThinker improves Mistral 7B instruct by 3.4%, LLaMA3 8B instruct (Touvron et al., 2023) by 6.5%, and Qwen 2 7B by 7% on two logical benchmarks and two mathematics benchmarks. We further demonstrate that TypedThinker can directly be applied to the benchmark Contexthub (Hua et al., 2024) and outperforms other baselines. Moreover, we show that the meta-thinker can be adapted in much larger LLMs such as GPT-4o or domain-specific LLMs such as MetaMath (Yu et al., 2023) and enhance their reasoning.
@@ 2 Related Work
Logical Reasoning Logical reasoning includes various methods to emulate human-like thought processes (Wason and Johnson-Laird, 1972; Dowden, 2018; Nunes, 2012). Deductive reasoning focuses on deriving specific conclusions from general principles or premises, ensuring that conclusions logically follow if the premises are true (Johnson-Laird, 2010). In contrast, inductive reasoning involves generalizing from specific instances to broader principles, often used to identify patterns and make predictions based on empirical data (Flach and Kakas, 2000). Abductive reasoning, considered more creative and open-ended, involves forming hypotheses to explain observations, often generating the most plausible explanation rather than a guaranteed conclusion (Douven, 2011). Analogical reasoning is concerned with the comparison between two or more objects and drawing a conclusion based on the similarity (Bartha, 2013). Previous LLMs studies on logical reasoning mainly focus on benchmarking its performance in different reasoning types (Bang et al., 2023; Dougrez-Lewis et al., 2024; Luo et al., 2023; Yu et al., 2024), or applying one reasoning type to solve the corresponding reasoning problems, such as using inductive reasoning for inductive reasoning problems (Wang et al., 2023a; Shao et al., 2024; Yang et al., 2024). Instead, this paper mainly focuses on the selection and application of the appropriate reasoning type when solving a general logic or math problem.
@@
Reasoning in Large Language Models Plenty of studies have been done to enhance the reasoning capability of LLMs. Chain-of-thoughts methods focus on creating better instructions to improve the quality of the reasoning process, such as Complex CoT (Fu et al., 2023), Tree of Thought (ToT) (Yao et al., 2023) and Graph of Thought (Besta et al., 2024). Refinement-based methods revise LLMs solutions by the feedback from themselves or others model (Akyürek et al., 2023; Wang and Li, 2023). Search-based methods use the reward model to search the best reasoning path (Lightman et al., ; Liu et al., 2023b; Hao et al., 2023). While most focus on creating high-quality reasoning paths, the diversity of thinking attracted more attention recently. Studies have investigated the diversity brought by repeated sampling (Brown et al., 2024) or multi-agent discussion with different prompts (Du et al., 2023; Liang et al., 2023; Suzgun and Kalai, 2024). Our paper aims to diversify thinking by incorporating suitable reasoning types for each instance.
@@
Self-improvement and Self-training in LLMs Recent works explore the self-improvement capability of LLMs, by finetuning LLMs on their high-quality generations (Wang et al., 2023b; Huang et al., 2023; Toshniwal et al., 2024). This process can be extended to multiple iterations Gülçehre et al. (2023); Aksitov et al. . Benefiting the LLMs’ ability to follow instructions, researchers also ask LLMs to provide feedback themselves and improve their responses without finetuning (Peng et al., 2023; Shinn et al., 2023). This can be further enhanced by using their own feedback as the reward model to provide better signals for finetuning (Yuan et al., 2024; Kumar et al., 2024). In this paper, we focus on stimulating their capabilities to conduct various reasoning types and use these experiences to diversify their thinking in reasoning type selection and following.
@@ 3 TypedThinker: Diversify Thinking with Typed Reasoning
In this paper, we focus on four logical reasoning types: deductive, inductive, abductive, and analogical reasoning defined in  (Nunes, 2012). For each reasoning type, we provide a short definition and a simple example to demonstrate the inference rules, which are listed in Table 8 in the Appendix. Based on that, we introduce a reasoning framework TypedThinker to diversify LLMs’ thinking with different reasoning types. As shown in Figure 2, there are three components in TypedThinker: the meta-thinker to select reasoning type, explicit collection for demonstration, and the LLM reasoner to exploit one particular reasoning type. TypedThinker optimizes the implicit policy of the meta-thinker and updates the explicit collection of demonstration based on previous experiences.
Figure 2: TypedThinker consists of three components: the meta-thinker to select the reasoning types, the explicit collection of demonstrations to retrieve relevant experience, and the reasoner to conduct the specific reasoning. The meta-thinker is fine-tuned to predict an effective score $s\in[0,1]$ for each reasoning type.
@@
3.1 Typing Reasoning with Implicit Policy and Explicit demonstration
@@
Let $D=\{(\bm{x_{1}},y_{1}),\cdots,(\bm{x_{N}},y_{N})\}$ be a set of $N$ problems, where $\bm{x}_{i}$ and $y_{i}$ is the problem and the ground-truth answer of the $i-$th instance. We define a reasoning type space $\mathcal{F}$ that includes an empty type and four types of reasoning: deductive, inductive, abductive, and analogical. The goal is to model the selection and implementation of various reasoning types as implicit and explicit collection of demonstration, thus enhancing LLMs’ performance in reasoning tasks.
@@
Meta-thinker for the reasoning type identification. Given a problem $\bm{x}$, the goal of the meta-thinker is to select an appropriate set of reasoning types to solve the problem. Specifically, it predicts an effectiveness score $s_{k}\in[0,1]$ for each reasoning type $f_{k}\in\mathcal{F}$, which can be represented as $s_{x,k}=\pi_{\theta}(\bm{x},f_{k})$. $s_{x,k}=0$ indicates that the problem $\bm{x}$ can hardly be solved by the reasoning type $f_{k}$ with a limited sampling times  [^2] In this paper, we sample at most 10 times for one problem.. Note that the effectiveness scores of different types are independent of each other and their sums are not necessary to be 1. The most effective reasoning type is defined as $f^{*}(\bm{x})=\arg\max_{f_{k}\in\mathcal{F}}s_{x,k}$. Meanwhile, we can obtain a set of reasoning types with a non-zero effective ratio, which we call the effective set: $F(\bm{x})=\{f_{k}|s_{x,k}>0\}$. We initialize $\pi_{\theta}$ with a pre-trained LLM. The prompt is listed in Appendix A.1.
@@
Explicit collection of demonstration TypedThinker collects a set of demonstration $M=\bigcup_{f_{k}\in\mathcal{F}}M_{k}$ for each type of reasoning. For each problem in the training set, we keep one correct solution per reasoning type if applicable, resulting in a set of at most $|D|\times|\mathcal{F}|$ solutions. If multiple solutions exist for one problem, we keep the longest to get a more detailed context. The entry in the collection of demonstration $\mathcal{M}_{k}$ is represented as a tuple of $(\bm{x}_{r}^{k},sol_{r}^{k})$. Here $sol_{r}^{k}$ is the concrete reasoning process of the reasoning type $f_{k}$, including the predicted answer. During inference stage, given a new problem $\bm{x}$ and its reasoning type $f_{k}$, it retrieves a set of relevant experience $d_{x}^{k}=\{(\bm{x}_{r}^{k},sol_{r}^{k})\in M_{k}|L(\bm{x}_{r}^{k},\bm{x})<\delta\}$. $L$ is the distance function measuring relevancy between two problems and $\delta\in[0,1]$ is the relevancy threshold. We use the cosine similarity between the semantic embeddings as the distance function. The retrieved experiences are used as the few-shot examples of the reasoner.
@@
Reasoner to perform the reasoning according to the type. The reasoner applies the reasoning type $f_{k}$ to the problem $\bm{x}$ and provides a detailed reasoning path for its predicted answer $\hat{y}$. The reasoner is based on LLM to conduct reasoning and the instruction is composed of $(\bm{x},f_{k},d^{k})$, where the $d^{k}$ is the retrieved relevant successful experience. The reasoner can be further optimized via instruction tuning to enhance the capability of conducting a specific type of reasoning.
@@
To conclude, TypedThinker enhances LLMs’ reasoning by estimating the effectiveness of each reasoning type and guiding the LLM reasoner with the demonstration of this reasoning type. Specifically, the meta-thinker $\pi_{\theta}$ predicts an effective score $s_{k}$ for each reasoning type. TypedThinker then retrieves the most relevant reasoning demonstration $d^{k}$ from the fixed explicit collection corresponding to the reasoning type $f_{k}$. Finally, the reasoners conduct the specific type of reasoning $f_{k}$ with the guidance of demonstration. In our experiments, we use two approaches to decide the reasoning types used in the problem-solving: One is to greedily resample several times based on the most effective reasoning type $f^{*}$ and use self-consistency (Wang et al., 2022) to enhance the answer; the other is to sample solutions for all effective reasoning types $F$, and apply a weighted vote with the effective score as the coefficient. By default, we use the greedy approach for TypedThinker, and we discuss the weighted vote in Section 4.
3.2 Optimize Implicit Policy for Reasoning Type Selection and Exploiting